OpenStudy (ximstorm):

Which of these shows 7 + 2y rewritten using the commutative property of addition? 7y + 2 2y + 7 7 − 2y 2y − 7

jimthompson5910 (jim_thompson5910):

the `commutative property of addition` is a fancy way of saying "you can add 2 numbers in any order" examples: 1+2 = 3 2+1 = 3 7+8 = 15 8+7 = 15 29+11 = 40 11+29 = 40 and so on...

jimthompson5910 (jim_thompson5910):

This applies to algebra because variables are just placeholders for numbers, so something like x+y is the same as y+x

OpenStudy (ximstorm):

B!

OpenStudy (anthonyym):

correct

jimthompson5910 (jim_thompson5910):

very good @ximstorm